OpenStudy (anonymous):

What is the point-slope form of an equation with the slope of 3/5 that passes through the point (10, –2)? A. y + 10 = 3/5(x – 2) B. y + 2 = 3/5(x – 10) C. y – 10 = 3/5(x + 2) D. y – 2 = 3/5(x + 10)

OpenStudy (anonymous):

Those cup size is the answer no lie

OpenStudy (anonymous):

lmao wow umm thanks I guess? this isn't for me anyways just helping a friend out with her math pre-test

OpenStudy (anonymous):

The equation for point slope form is \[y-y _{1}=m(x-x _{1})\] So substitute the equation with your given values Remember that m is the slope and x1 is the x coordinate, y1 is the y coordinate

OpenStudy (anonymous):

So, your point-slope form would be: y-(-2)=3/5(x-10), which would simplify to y+2=3/5(x-10)

OpenStudy (anonymous):

So, the answer is B., y+2=3/5(x-10) @jessica_the_meow

OpenStudy (anonymous):

ty for explaining :) your were very helpful :)

OpenStudy (anonymous):

Np :)

OpenStudy (anonymous):

And for future reference, if you are ever given two points rather than one for this kind of problem, remember that you are only paying attention the first point (from left to right), because you only need x1 and y1, not x2 and y1 for point-slope form

OpenStudy (anonymous):

..Happy studying! :)

OpenStudy (anonymous):

okay ill keep that in mind :) @openstudier101x

OpenStudy (anonymous):

..er sorry typo haha not x2 and y2* for point-slope form

OpenStudy (anonymous):

..But great! Glad to help :)